OpenStudy (4meisu):

a. 38.0 mL of 0.10M HCl is needed to neutralize 25.0 mL to NH4OH. What is the molarity of the ammonium hydroxide?

OpenStudy (anonymous):

I would first write out a balanced equation for the reaction

OpenStudy (anonymous):

HCl + NH4OH ---> H20 + NH4Cl, Now you can use some stoichometry to figure out the moles of NH4oh that needs to be nuetralized.

OpenStudy (anonymous):

\[38.0 ml HCl (\frac{.10 mole HCl}{1000 ml HCL})(\frac{1 mole NH_4OH}{1 mole HCL})=.0038 mole NH_4OH\] Now we divide it by .025 L to get the molartity \[\frac{.0038 moles}{.025 L}=.152 M \]